# Think-to-Talk or Talk-to-Think? When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning

### v3

@@ 6 Conclusions
We conducted causal probing analyses of when (sub-)answers are determined in the CoT process, using synthetic arithmetic problems as a controlled testbed. Across a range of models and task difficulties, we found that models predominantly operated in a Talk-to-Think mode: they resolved the necessary subproblems during CoT generation phase. Moreover, causal intervention experiments revealed a strong recency bias linking hidden states to outputs, indicating that LLMs rely heavily on recent computations when generating explanations. This pattern further suggests that their internal reasoning flow largely aligns with the produced explanations.
@@
Comprehensiveness of experimental settings
Some experiments were conducted with a limited scope; for example, the experiments with various models in §˜4.3 are conducted only on the Level 3 task. Additionally, causal interventions (§˜5) are performed only with Qwen2.5-7B. Conducting our experiment with more models and tasks will further enhance the generalizability of the results.
Variety of task
We analyzed the internal reasoning patterns of language models using synthetic arithmetic reasoning tasks. The use of synthetic data allows for more detailed control compared to experiments on natural language tasks. However, vocabulary and expression diversity, for example, are limited compared to natural language tasks. Therefore, conducting similar analyses on reasoning tasks will verify whether the results of this study apply to other broader, realistic contexts as well. Additionally, in our study, we focus on a single reasoning-chain pattern, and it would be desirable to also conduct experiments using other reasoning chain strategies and formats. On the other hand, controlling the length and granularity of them is difficult because there are various options. Conducting experiments for other reasoning chain strategies and formats is expected to provide more general insights.
@@
Interpreting internal mechanisms of LMs using probing have been actively conducted in our field Conneau et al. (2018); Tenney et al. (2019); Campbell et al. (2023); Li et al. (2023); however, there are criticisms regarding the validity of some probing approaches Liu et al. (2023); Burns et al. (2023). One way to overcome such concerns will be to analyze the generality of obtained results through more diverse methodologies Gurnee et al. (2023); Bricken et al. (2023).
@@
Appendix A Supplemental results
Model Level Task
Qwen2.5 (7B) 1 100
2 100
3 100
4 100
5 100
Qwen2.5 (14B) 3 100
Qwen2.5 (32B) 3 100
Qwen2.5-Math (7B) 3 100
Yi1.5 (9B) 3 100
Yi1.5 (34B) 3 100
Llama3.1 (8B) 3 100
Llama3.2 (3B) 3 97.6
Mistral-Nemo (12B) 3 99.6
@@ Appendix A Supplemental results
A.1 Performance of language models in the arithmetic tasks
@@
A.2 All probing results
Figures 9 through 53 present the probing results for all models and tasks discussed in this paper. Tables 6 through 20 summarize these results for thresholds ($\tau$) ranging from 0.85 to 0.95.
From these results, we observe trends similar to those described in §˜4.3 across many settings. However, for the smaller model Llama3.2 (3B), increasing the threshold $\tau$ often leads to cases where the accuracy does not reach the threshold (N/A). Nonetheless, a consistent pattern remains: $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low whereas $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high, indicating a Talk-to-Think mode.
Appendix B Hyperparameters
Table 5 shows the hyperparameters used for training the probes.
@@ Appendix B Hyperparameters
Appendix C Addtional causal intervention results
Figures 7 and 8 show the causal intervention results for the target tokens $z_{17}$ and $z_{32}$, respectively. Here, in addition to the Success rate, we also present the Unchanged rate as a metric. The Unchanged rate indicates how frequently (%) the intervened output $\hat{\tilde{y}}$ remains the same as $y$. If this value is small, it indicates that the patched hidden states do not affect the output.
@@ Appendix E Usage of AI assistants
Table 6: Results for various models on the task Level 1 ($\tau=0.85$).
@@
Table 7: Results for various models on the task Level 2 ($\tau=0.85$).
@@
Table 8: Results for various models on the task Level 3 ($\tau=0.85$).
@@
Table 9: Results for various models on the task Level 4 ($\tau=0.85$).
@@
Table 10: Results for various models on the task Level 5 ($\tau=0.85$).
@@
Table 11: Results for various models on the task Level 1 ($\tau=0.90$).
@@
Table 12: Results for various models on the task Level 2 ($\tau=0.90$).
@@
Table 13: Results for various models on the task Level 3 ($\tau=0.90$).
@@
Table 14: Results for various models on the task Level 4 ($\tau=0.90$).
@@
Table 15: Results for various models on the task Level 5 ($\tau=0.90$).
@@
Table 16: Results for various models on the task Level 1 ($\tau=0.95$).
@@
Table 17: Results for various models on the task Level 2 ($\tau=0.95$).
@@
Table 18: Results for various models on the task Level 3 ($\tau=0.95$).
@@
Table 19: Results for various models on the task Level 4 ($\tau=0.95$).
@@
Table 20: Results for various models on the task Level 5 ($\tau=0.95$).

### v4

@@ 6 Conclusions
We conducted probing experiments to determine when (sub)answers are derived during the CoT-style reasoning, using synthetic arithmetic problems as a controlled testbed. Across multiple models and task difficulties, we found that models tend to resolve the necessary (sub)answer after the CoT begins, i.e., computing the answer on the fly during CoT generation. Moreover, causal experiments support that the intervention to the CoT part specifically and causally impacts the final answer; that is, we conclude that the CoT is faithful to the final answer at least in our controlled experimental settings.
@@
Variety of tasks
We analyzed the internal reasoning patterns of language models using synthetic arithmetic reasoning tasks. The use of synthetic data allows for more detailed control compared to experiments on natural language tasks. However, vocabulary and expression diversity, for example, are limited compared to natural language tasks. Therefore, conducting similar analyses on more realistic reasoning tasks in natural language will verify whether the results of this study apply to other broader, realistic contexts as well. In addition, we focused on the tasks that presumably require step-by-step reasoning as the CoT process, but as suggested in existing studies (§˜5), the situation may be different if the task is, in the first place, so simple that CoT is no longer needed.
@@
Interpreting internal mechanisms of LLMs using probing has been actively conducted in our field Conneau et al. (2018); Tenney et al. (2019); Campbell et al. (2023); Li et al. (2023); however, there are criticisms regarding the validity of some probing approaches Liu et al. (2023); Burns et al. (2023). One way to overcome such concerns will be to analyze the generality of obtained results through more diverse methodologies Gurnee et al. (2023); Bricken et al. (2023).
Causal intervention purity
Hidden states carry mixed information (e.g., features of the input text itself). Thus, even though activation patching is standard in prior work, we cannot fully rule out noise introduced by this intervention.
@@
Acknowledgments
This work was supported by JST CREST Grant Number JPMJCR20D2, JSPS KAKENHI Grant Numbers JP25KJ0615 and JP25K03175, JST SPRING Grant Number JPMJSP2114, Google Research grant, and the Nakajima Foundation. Ana Brassard’s contribution was supported by a RIKEN Incentive Research Project (FY2024). Part of this work was carried out using the computer resource offered under the category of “General Projects” by Research Institute for Information Technology, Kyushu University. We thank the member of the Tohoku NLP Group for their cooperation in this research.
Appendix A Alternative CoT formats.
For comparison, we also ran probing experiments on Qwen2.5-7B under two reasoning chain formats. Specifically, we defined two formats. The first is the Simple CoT setting, in which the reasoning chain outputs only the intermediate sub-results. The second is the Implicit (reasoning) setting, which omits intermediate computation steps in the output text. Table 5 lists example equations.
The experimental results are shown in Table 6 and Figures 57 and 58. Under the Simple CoT setting, the model’s task accuracy was 99.5%, whereas under the Implicit reasoning setting it was 77.8% [^6] Please note that, unlike the other experimental configurations, under the implicit reasoning setting the model is unable to answer the problems with sufficiently high accuracy.. As in the Simple CoT setting, intermediate results are linearly separable in the Output segment ($t^{*}_{\mathrm{eq}}>0$). This trend is consistent with the results obtained using the default CoT format in §˜3.3. In the implicit reasoning setting, the accuracy did not reach the threshold at any position (both $\mathrm{Acc}_{\mathrm{\prec CoT}}$ and $\mathrm{Acc}_{\mathrm{\succ CoT}}$ are N/A). These results suggest that when the model fails to solve a problem, the corresponding (sub-)answers are also unlikely to be represented in its internal states. This, in turn, indicates that the answers decoded from internal states are faithful, in the sense that when the model cannot solve the problem, there may be no linearly decodable representation either.
Appendix B Supplemental results
Level
1 2 3 4 5
Qwen2.5 (7B) 100.00 100.00 100.00 100.00 100.00
Qwen2.5 (14B) 100.00 100.00 100.00 100.00 98.25
Qwen2.5 (32B) 100.00 100.00 100.00 100.00 100.00
Qwen2.5-Math (7B) 100.00 100.00 100.00 100.00 100.00
Yi1.5 (9B) 100.00 100.00 100.00 100.00 100.00
Yi1.5 (34B) 100.00 100.00 100.00 100.00 100.00
Llama3.1 (8B) 100.00 100.00 100.00 100.00 99.40
Llama3.2 (3B) 99.95 99.75 93.15 90.90 38.45
Mistral-Nemo (12B) 100.00 100.00 99.95 100.00 99.85
@@ Appendix B Supplemental results
B.1 Model performance on arithmetic tasks
@@
B.2 Additional probing predictions for correct/incorrect instances.
Figure 10 shows the top-1 predictions of the probe for $\mathrm{B}$ in instances where Llama3.2-3B produced incorrect answers for Task 3. In contrast, Figure 11 illustrates the probe predictions for instances where Llama3.2-3B generates correct responses. In Figure 3 the model should output (B=4) but instead produces 0. We observe the same pattern in other examples in Figure 10. One possible cause is that digits for computed results (e.g., 4) and digits from the input (e.g., 0) are represented in neighboring subspaces, which may cause confusion; we leave a detailed analysis for future work.
@@
B.3 All probing results
Figures 12 through 56 present the probing results for all models and tasks discussed in this paper. For each figure, the input sequence below the graphs is one example (the results are averaged over the test set). The upper part indicates the maximum probing accuracy achieved at each token position $t$. The bottom part shows the probing accuracies in each token $t$, layer $l$, and variable $v_{i}$. Tables 9 through 23 summarize these results for thresholds ($\tau$) ranging from 0.85 to 0.95. From these results, we observe trends similar to those described in §˜3.3 across many settings. However, for the smaller model Llama3.2 (3B), increasing the threshold $\tau$ often leads to cases where the accuracy does not reach the threshold (N/A). Nonetheless, a consistent pattern remains: $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low whereas $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high.
B.4 Additional causal intervention results
Figures 59111 show the causal intervention results for the target tokens $z_{17}$ and $z_{32}$, respectively.
Here, in addition to the Success rate, we also present the Unchanged rate as a metric. The Unchanged rate indicates how frequently (%) the intervened output $\hat{y}^{\mathrm{patch}}$ remains the same as $y$. If this value is small, it indicates that the patched hidden states do not affect the output.
Additional discussion: Memory vs. recomputation.
Focusing on the case where the target token is ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}\textbf{5}}}_{\hskip 1.42271pt\mathbf{4}}$ in Figure 6, if the model were recomputing when generating this 5, it would be expected to show a causal relationship with the segment ${\underline{{\mathrm{B}=2+3}}}_{\hskip 1.42271pt\mathbf{1}}$, where the pre-computation equation information is expected to be explicitly represented. However, in practice, a strong causal relationship was observed only with the hidden state of the immediately preceding segment ${\underline{\mathrm{B}=5}}_{\hskip 1.42271pt\mathbf{2}}$. This pattern suggests that the model is not recomputing; instead, it likely relies on the immediately preceding intermediate result stored in the CoT text.
Appendix C Hyperparameters
Table 8 shows the hyperparameters used for training the probes.
@@ Appendix E Usage of AI assistants
Figure 10: Error analysis of cases where Llama3.2-3B generated incorrect answers. The vertical axis represents the index of the transformer layer. The horizontal axis represents the tokens input to the model over time. The numbers in the figure indicate the labels predicted by each probe (top-1) at each layer and time step. The numbers highlighted in green represent the gold labels for the predictions, while those highlighted in red denote the values incorrectly generated by the model.
Figure 11: Top 1 predictions of probes where Llama3.2-3B generated correct answers. The vertical axis represents the index of the transformer layer. The horizontal axis represents the tokens input to the model over time. The numbers in the figure indicate the labels predicted by each probe (top-1) at each layer and time step. The numbers highlighted in green represent the gold labels for the predictions.
@@
Table 9: Results for various models on the Level 1 task ($\tau=0.85$).
@@
Table 10: Results for various models on the Level 2 task ($\tau=0.85$).
@@
Table 11: Results for various models on the Level 3 task ($\tau=0.85$).
@@
Table 12: Results for various models on the Level 4 task ($\tau=0.85$).
@@
Table 13: Results for various models on the Level 5 task ($\tau=0.85$).
@@
Table 14: Results for various models on the Level 1 task ($\tau=0.90$).
@@
Table 15: Results for various models on the Level 2 task ($\tau=0.90$).
@@
Table 16: Results for various models on the Level 3 task ($\tau=0.90$).
@@
Table 17: Results for various models on the Level 4 task ($\tau=0.90$).
@@
Table 18: Results for various models on the Level 5 task ($\tau=0.90$).
@@
Table 19: Results for various models on the Level 1 task ($\tau=0.95$).
@@
Table 20: Results for various models on the Level 2 task ($\tau=0.95$).
@@
Table 21: Results for various models on the Level 3 task ($\tau=0.95$).
@@
Table 22: Results for various models on the Level 4 task ($\tau=0.95$).
@@
Table 23: Results for various models on the Level 5 task ($\tau=0.95$).
@@
Figure 57: Probing results when Qwen2.5-7B solves Level 2 with Simple CoT.
Figure 58: Probing results when Qwen2.5-7B solves Level 2 with Implicit reasoning.
Figure 59: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 60: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 61: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 62: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 63: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 64: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 65: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 66: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 67: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 68: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 69: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 70: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 71: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 72: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 73: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 74: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 75: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 76: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 77: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 78: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 79: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 80: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 81: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 82: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the success rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 83: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the success rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 84: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 85: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 86: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 87: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 88: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 89: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 90: Results of the causal intervention on Qwen2.5-14B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 91: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 92: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 93: Results of the causal intervention on Qwen2.5-32B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 94: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 95: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 96: Results of the causal intervention on Qwen2.5-Math-7B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 97: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 98: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 99: Results of the causal intervention on Yi-1.5-9B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 100: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 101: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 102: Results of the causal intervention on Yi-1.5-34B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 103: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 104: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 105: Results of the causal intervention on Llama-3.1-8B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 106: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 107: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 108: Results of the causal intervention on Llama-3.2-3B. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 109: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the unchanged rate when the intermediate token $z_{17}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}})$ is the target token.
Figure 110: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the unchanged rate when $z_{32}$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}})$ is the target token.
Figure 111: Results of the causal intervention on Mistral-Nemo-Base-2407. Each grid cell shows the unchanged rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.